# Data-Model Co-Evolution: Growing Test Sets to Refine LLM Behavior

### v1

@@ 3.2.3. Test Sets are Diverse, Application-Specific, and Often Ad-Hoc
  • Out-of-Scope Inputs: For customer-facing systems (e.g., banking chatbots), a recurring concern was how the model handles irrelevant or nonsensical queries. Participants noted that real users often “try anything” either out of curiosity about the AI or to probe its boundaries. Tests in this category therefore evaluate both detection (is the request off-domain?) and response style (is the refusal informative and consistent).
@@ 3.2.4. Evaluation Dichotomy Between Rapid Gut-Checks and Costly Metrics
For the frequent, small-scale iterations typical of prompt engineering, the dominant method is the “gut-check”—“In the end, what really mattered was more of a gut feeling—like, ‘this version feels more concrete and actually covers the things we need.”’ [P4]. Practitioners make a small change and then manually observe a handful of outputs to get a “feel” for the impact. This approach is valued for its speed. As one participant described, “Three developers ran one–two-hour jam sessions, throwing queries and checking logs. Roughly  100 queries in total; if results looked good, we locked in the system prompt.” [P3]. This informal, qualitative assessment is the engine of the day-to-day development loop.
However, to gain higher confidence or to justify changes to stakeholders, this intuitive sense is not enough. In these situations, teams turn to a more formal process: curating a test set, manually labeling each output, and calculating quantitative metrics. Providing the ”hard numbers” needed to demonstrate progress to clients or leadership.
This creates a central tension in their workflow. While practitioners desire the rigor of quantitative evaluation, its high cost in time and labor makes it impractical for every iteration. As one participant put it, “Human review by the representatives wasn’t feasible on every cycle. Practically, we could only do it about once a month.” [P4]. This dichotomy forces teams into a difficult compromise: they perform the majority of their iterations using fast but low-confidence gut-checks, punctuated by infrequent, high-effort formal evaluations. This leaves them without a continuous, clear signal of whether their iterative changes are truly improving the system.
@@ 3.3. Design Implications
Our formative study revealed a clear gap: while practitioners are locked in a rapid, iterative loop of prompt refinement, they lack systematic tools to support this process. These findings motivated the design of a workflow grounded in a new paradigm.
@@
Toward a New Paradigm. Synthesizing these implications, we identified a fundamental disconnect in current workflows: the prompt specification (model) and the test user inputs (data) are treated as separate, informally managed artifacts. A change to one does not systematically inform or validate the other. This insight led us to propose a new paradigm for LLM development: data–model co-evolution. In this paradigm, the test set and the specification are not independent but are treated as intertwined artifacts that must grow and evolve in tandem. Each discovered failure (data) must durably impact the specification (model), and each change to the specification must be validated against all accumulated data. Our system is a direct embodiment of this paradigm.
Figure 2. System interface for the Co-Evolution Workflow. The left panel supports authoring and tracking prompt instructions across versions. The right panel manages data: the top section maintains a living test set with model responses from both prior and current prompt instruction versions, while the bottom section displays generated candidate inputs for user review. Together, these panels embody the co-evolution principle by tightly coupling specification revisions with test set examples. The examples in this figure, originally written in Korean, were used in the user study and translated into English.
@@ 4. Interactive Co-Evolution Workflow
To operationalize our paradigm of data–model co-evolution, we designed and built an interactive system for refining LLM behavior. We first describe the design principles that guided the workflow (4.1), then introduce the interface (4.2), and finally explain the workflow in detail (4.3).
@@ 4.1. Design Principles
Before presenting the principles, we first clarify our terminology. In this work. In the context of LLM-based applications, we use the term model in a broadened sense. Specifically, the model refers to the prompt instruction (a.k.a., system prompt) that encodes the intended policy, which is concatenated with a given user input and sent to the LLM API to produce a response. By contrast, we use the term data to denote end users’ inputs which invoke the model to produce different outputs. In other words, the model remains stable as the governing specification, while the data varies with every user interaction. With this in place, we describe our key design principles:
4.1.1. From Model to Data: Using the Specification to Discover Its Own Blind Spots.
The refinement process begins with the current model (i.e., prompt instruction). Its weaknesses are revealed not in the average case, but in specific, unforeseen edge cases. We leverage the powerful data synthesis capabilities of modern LLMs to find these weaknesses. The system uses the current prompt instruction to generate challenging inputs, which surfaces potential failure modes and helps the user ideate and anticipate problematic cases.
4.1.2. From Data to Model: Grounding Specification Refinements in Concrete Evidence.
Once the edge cases (data) are surfaced, they become the foundation for refining the prompt instruction (model). This principle ensures that changes are not based on vague intuition but on concrete evidence. As a user confronts a generated failure, they are prompted to label it and articulate a rationale. This step can help clarify their own internal policy. It is then used to guide the update to the prompt instruction.
4.1.3. A Living Test Set:
Each discovered case is preserved in a growing test set that reflects the current state of the model. This evolving benchmark enables users to assess overall progress, detect regressions, and ensure systematic improvement.
@@ 4.1.4. LLM-Assisted, Human-Governed Workflow
4.1.5. Scaffolding Generalization: From a Single Case to a General Rule:
The workflow aims to support translating the insight from single failures into generalizable rules. Our workflow explicitly scaffolds this transition by allowing users to test that a rationale against similar cases, before committing it to the model update. This helps transform a reactive fix (”this one output was bad”) into a proactive policy enhancement (”this class of outputs is undesirable, and here is a rule to prevent it”), building users confidence and leading to more robust models.
@@ 4.2. System Interface
Figure 2 shows the interface. It consists of two main columns that physically separate the model specification from the data. The left panel is dedicated to editing and tracking the prompt instruction, while the right panel is for discovering and curating the test data.
4.2.1. The Prompt Instruction (Model) Panel
@@
Prompt Instruction Editor: At the top is a text editor where developers write and refine the instructions that guide the LLM’s behavior.
Version History: Below the editor, a history view lists all saved versions of the prompt instruction. It also serves as a performance dashboard. Each version is visualized with a percentage bar chart to show the performance metric of its overall accuracy calculated against the current test set. This allows users to see at a glance whether a revision improved performance or caused an issue. Users can easily revert to or compare against previous versions.
@@ 4.2.2. The Data Panel: Discovering and Accumulating Evidence
The Test Set: The main component is the table representing the test set. It is a collection of labeled examples that serve as the ground truth for the application’s policy. Its columns display user inputs, model responses, Good/Bad ratings, and the corresponding rationales. When a new prompt version is evaluated, this table updates to show responses from both the old and new versions side-by-side, making the impact of any change immediately apparent.
Generated Examples: Below the test set, a smaller table serves as a staging area for examples synthesized by LLMs. This is where new, unlabeled edge cases are surfaced for the user to review. When an example from this area is labeled, it is formally promoted into the test set above.
@@ 4.3. Co-Evolution Workflow in Action
A user’s interaction with the system follows an iterative cycle designed to discover failures, generalize from them, and refine the model’s specification against the test set, which is illustrated in Figure 3. Below, we describe each step of the workflow, using the example of a banking customer-support chatbot that should only handle finance-related queries and politely redirect queries that are out of scope.
Step 1: Discovering Failures. The workflow begins in the Generated Examples panel, where the system generates candidate user inputs intended to be “hard” cases under the current prompt instruction. These examples are generated by an LLM with the explicit goal of probing boundaries that are likely ambiguous or underspecified. Users inspect the model’s responses to these inputs and labels them as Good or Bad. For example, the chatbot might receive the input “How should I manage my credit score while preparing for job applications?”. If the AI refuses to answer on the grounds that this is an “employment-related” question, the user may label this as Bad, reasoning that credit score management is central to their banking support even if the query also mentions job preparation. While some applications might choose to withhold an answer in such cases, our scenario reflects a policy preference to provide guidance as long as the query does not cross into problematic territory.
Step 2: Articulating a Rationale. When a user labels an example as Bad, the system prompts them to provide a rationale explaining why the behavior was incorrect. Instead of leaving this as a free-form burden, the interface provides two LLM-suggested options, but the users can also write their own. For example, they might enter “Questions about credit scores should be answered, even if they are mentioned in the context of employment or other life events.” This articulation transforms an implicit judgment into an explicit principle, shifting the mindset from reactive correction (“this answer is wrong”) to policy refinement (“here is the scope rule that the chatbot should follow”).
@@
Step 3: Generalizing the Rationale with Neighborhood Probing. The aim of this step is to move from a single failure to a robust, reusable rule. Our workflow supports neighborhood probing: given the user’s rationale, the system synthesizes a small set of semantically similar inputs that preserve the same triggers while varying context. Clicking a button on the rationale popup opens a larger popup for this purpose. As shown in Figure 4, users can see and label these cases to see if the rationale consistently predicts undesired behavior. If the model mishandles all of them, the rationale is confirmed as broadly relevant; if not, the developer may refine the rationale to be more specific. These examples will be added to the test set. We crafted a prompt that generates three new examples as below:
Generate three new user inputs to verify if the failure described in the Human Rationale is reproducible,
without widening scope.
@@
- Human Rationale (root cause & semantic triggers): {human_rationale}
@@
Step 5: Evaluating Against the Test Set. Upon applying the revision, several actions are triggered automatically. First, the Prompt Instruction panel is updated with the new version. Second, the original failed example and the additional examples from the neighborhood probe are added to the Test Set panel. The panel now runs the new prompt instruction against all test cases.
@@ 4.4. Implementation Details
Our system is implemented as a web-based application that runs directly in the browser. The user-facing interface is built with React, while components that require LLM calls are handled by a Python server developed with Flask. The server communicates with external LLM APIs to support various stages of the workflow.
@@
We primarily used Gemini 2.5 Flash (Comanici et al., 2025) from Google Gemini API [^1] https://ai.google.dev/gemini-api/docs/models to power the workflow steps, such as generating synthetic examples, retrieving similar examples, and suggesting prompt revisions. To ensure fast interactions, reasoning-heavy “thinking mode” was avoided in most cases, with the exception of generating similar examples. For model responses to end users, where prompt instructions are simply combined with user inputs, we employed lighter models such as Gemini 2.5 Flash Lite.

### v2

@@ 3.2.3. Test Sets are Diverse, Application-Specific, and Often Ad-Hoc
  • Out-of-Scope Inputs: For customer-facing systems (e.g., banking chatbots), a recurring concern was how the model handles irrelevant or nonsensical queries. Participants noted that real users often “try anything” either out of curiosity about the AI or to probe its boundaries. Tests in this category therefore evaluate both detection (is the request off-domain?) and response style (is the refusal informative and consistent?).
@@ 3.2.4. Evaluation Dichotomy Between Rapid Gut-Checks and Costly Metrics
For the frequent, small-scale iterations typical of prompt engineering, the dominant method is the “gut-check”: “In the end, what really mattered was more of a gut feeling. It’s like, this version feels more concrete and actually covers the things we need.” [P4]. Practitioners make a small change and then manually observe a handful of outputs to get a “feel” for the impact. This approach is valued for its speed. As one participant described, “Three developers ran one–two-hour jam sessions, throwing queries and checking logs. Roughly  100 queries in total; if results looked good, we locked in the system prompt.” [P3]. This informal, qualitative assessment is the engine of the day-to-day development loop.
However, to gain higher confidence or to justify changes to stakeholders, this intuitive sense is not enough. In these situations, teams turn to a more formal process: curating a test set, manually labeling each output, and calculating quantitative metrics. This provides the “hard numbers” needed to demonstrate progress to clients or leadership.
This creates a central tension in their workflow. While practitioners desire the rigor of quantitative evaluation, its high cost in time and labor makes it impractical for every iteration. As one participant put it, “Human review by the representatives wasn’t feasible on every cycle. Practically, we could only do it about once a month.” [P4]. This leaves teams with a difficult compromise: they perform the majority of their iterations using fast but low-confidence gut-checks, punctuated by infrequent, high-effort formal evaluations. This leaves them without a continuous, clear signal of whether their iterative changes are truly improving the system.
@@ 3.3. Design Implications
Our formative study revealed a clear gap: while practitioners are locked in a rapid, iterative loop of prompt refinement, they lack systematic tools to support this process. These findings motivated the design of a new workflow.
@@
Synthesizing these implications, we identified a fundamental disconnect in current workflows: prompt specifications and test inputs are treated as separate, informally managed artifacts. A change to one does not systematically inform or validate the other. This insight led us to propose a new integrated workflow for LLM development: data-prompt co-evolution, where the test set and the specification are not independent but are treated as intertwined artifacts that must grow and evolve in tandem. Each discovered failure must durably impact the specification, and each change to the specification must be validated against all accumulated data.
@@ 4. Interactive Co-Evolution Workflow
Figure 2. System interface for the Co-Evolution workflow. This example illustrates a content moderation scenario for a university student forum. The left panel supports authoring and tracking prompt instructions across versions. The right panel manages data: the top section maintains a living test set with model responses from both prior and current prompt instruction versions, while the bottom section displays generated candidate inputs for user review. Together, these panels embody the co-evolution principle by tightly coupling specification revisions with test set examples.
A screenshot of the system presents a two-column interface. The left side contains a prompt instruction editor, along with a version history panel that lists prompt revisions, each displaying percentage bar charts indicating performance. The right side manages data. The upper table, labeled Test Examples, lists user inputs, model responses and and correctness ratings for both the previous and new prompt versions. The lower table, labeled Generated Candidate Examples, displays generated inputs and their corresponding model responses.
To operationalize our data-prompt co-evolution workflow, we design and build an interactive system for refining LLM behavior. We describe the design principles (4.1), introduce the interface (4.2), and explain the workflow in detail (4.3).
@@ 4.1. Design Principles
Before presenting the principles, we first clarify our terminology. In typical LLM-based applications, the final input passed to an API is often constructed via a template that combines a static prompt instruction (a.k.a. system prompt) with dynamic user input (e.g., “Summarize the following text: {user_text_to_summarize}”). In the context of data-prompt co-evolution, we use the term prompt to refer specifically to the prompt instruction (or system prompt), as this serves as the editable specification encoding the intended policy. Conversely, we use the term data to denote the test set (or validation set) consisting of end user inputs which invoke the model to produce different outputs in response to them. With this in place, we describe our key design principles:
4.1.1. From Prompt to Data: Using the Specification to Discover Its Own Blind Spots.
The refinement process begins with the current prompt instruction. Its weaknesses are revealed not in the average case, but in specific, unforeseen edge cases. We leverage the powerful data synthesis capabilities of modern LLMs to find these weaknesses. The system uses the current prompt instruction to generate challenging inputs, which surfaces potential failure modes and helps the user ideate and anticipate problematic cases.
4.1.2. From Data to Prompt: Grounding Specification Refinements in Concrete Evidence.
Once the edge cases (data) are surfaced, they become the foundation for refining the prompt instruction. This principle ensures that changes are not based on vague intuition but on concrete evidence. As a user confronts a generated failure, they are prompted to label it and articulate a rationale. This step can help clarify their own internal policy. It is then used to guide the update to the prompt instruction.
4.1.3. A Living Test Set.
Each discovered case is preserved in a growing test set that reflects the current state of the prompt. This evolving benchmark enables users to assess overall progress, detect regressions, and ensure systematic improvement.
@@ 4.1.4. LLM-Assisted, Human-Governed Workflow
4.1.5. Scaffolding Generalization: From a Single Case to a General Rule.
The workflow aims to support translating the insight from single failures into generalizable rules. Our workflow explicitly scaffolds this transition by allowing users to test that a rationale against similar cases, before committing it to the update of prompts. This helps transform a reactive fix (“this one output was incorrect”) into a proactive policy enhancement (“this class of outputs is undesirable, and here is a rule to prevent it”), building users’ confidence and leading to more robust results.
The workflow consists of five steps. (1) Discovering Failures: the system synthesizes candidate inputs and runs them through the model. (2) Articulating a Rationale: the human labels the correctness of model responses; for incorrect cases, they write a short hypothesis explaining the error. (3) Probing with Neighborhoods: the system generates similar inputs around that rationale, and selected examples are added to the test set. (4) Refining the Specification: the prompt instruction is revised, for example by adding a new line, producing an updated version. (5) Evaluating against the Living Test Set: both prior and updated prompts are evaluated on the expanded test set, with outcomes marked as incorrect or correct.
@@ 4.2. System Interface
Figure 2 shows the interface. It consists of two main columns that physically separate the prompt specification from the data. The left panel is dedicated to editing and tracking the prompt instruction, while the right panel is for discovering and curating the test data.
4.2.1. The Prompt Instruction Panel
@@
Prompt Instruction Editor: At the top is a text editor where users write and refine the instructions that guide the LLM’s behavior.
Version History: Below the editor, the history panel lists all saved versions of the prompt instruction. It also serves as a performance dashboard. Each version is visualized with a percentage bar chart to show the performance metric of its overall accuracy calculated against the current test set. Green bars indicate the proportion of correct responses, while red bars indicate incorrect responses. Darker shades correspond to user-annotated labels, whereas lighter shades correspond to labels produced by LLM-as-judges (described further in Step 5 in Section 4.3). This allows users to see at a glance whether a revision improved performance or caused an issue. Users can easily revert to or compare against previous versions.
@@ 4.2.2. The Data Panel: Discovering and Accumulating Evidence
The Test Set: The main component is the table representing the test set. It is a collection of labeled examples that serve as the ground truth for the application’s policy. Its columns display user inputs, model responses, buttons where the user can indicate whether they think the responses are correct or incorrect, and the corresponding rationales. When a new prompt version is evaluated, this table updates to show responses from both the old and new versions side-by-side, making the impact of any change immediately apparent.
Generated Examples: Below the test set, another table serves as a staging area for examples synthesized by LLMs. This is where new, unlabeled edge cases are surfaced for the user to review. When an example from this area is labeled, it is formally promoted into the test set above.
@@ 4.3. Co-Evolution Workflow in Action
A user’s interaction with the system follows an iterative cycle designed to discover failures, generalize from them, and refine the specification against the test set, which is illustrated in Figure 3. Below, we describe each step of the workflow, using the example of a banking customer-support chatbot that should only handle finance-related queries and politely redirect queries that are out of scope.
Step 1: Discovering Failures. The workflow begins in the Generated Examples view, where the system generates candidate user inputs intended to be “hard” cases under the current prompt instruction. These examples are generated by an LLM with the explicit goal of probing boundaries that are likely ambiguous or underspecified. Users inspect the model’s responses to these inputs and labels them as Correct or Incorrect. For example, the chatbot might receive the input “How should I manage my credit score while preparing for job applications?”. If the AI refuses to answer on the grounds that this is an “employment-related” question, the user may label this as Incorrect, reasoning that credit score management is central to their banking support even if the query also mentions job preparation. While some applications might choose to withhold an answer in such cases, our scenario reflects a policy preference to provide guidance as long as the query does not cross into problematic territory.
Step 2: Articulating a Rationale. When a user labels an example as Incorrect, the system prompts them to provide a rationale explaining why the behavior was incorrect. Instead of leaving this as a free-form burden, the interface provides two LLM-suggested options, but the users can also write their own. For example, they might enter “Questions about credit scores should be answered, even if they are mentioned in the context of employment or other life events.” This articulation transforms an implicit judgment into an explicit principle, shifting the mindset from reactive correction (“this answer is wrong”) to policy refinement (“here is the scope rule that the chatbot should follow”).
@@
A modal window titled “Refine Prompt Instructions” demonstrates the neighborhood probing process. The interface displays a table where the first row, highlighted in yellow and labeled Original, shows a model failure case with a user-provided rationale for an incorrect rating. Below this row, three similar inputs synthesized by the system are listed. Each row includes a user input, the AI Response, and buttons for marking the response as correct or incorrect. Checkboxes on the left allow the user to select specific cases to be added to the test set. At the bottom, buttons labeled “Generate More Examples” and “Next Step” are displayed.
Step 3: Generalizing the Rationale with Neighborhood Probing. The goal of this step is to generalize a single failure into a reusable rule. Our workflow supports neighborhood probing: given the user’s rationale, the system synthesizes a set of semantically similar inputs that preserve the same triggers while varying context. Clicking a button on the rationale popup opens a larger popup for this purpose. As shown in Figure 4, users can label these cases to assess whether the rationale consistently predicts undesired behavior. If the model mishandles all of them, the rationale is confirmed as broadly relevant; if not, users may refine the rationale to be more specific. These examples will be added to the test set. We crafted a prompt that generates three new examples as follows:
Generate three new user inputs to verify if the failure
described in the Human Rationale is reproducible, without
widening scope.
@@
- Human Rationale (root cause & semantic triggers): {rationale}
@@
Step 5: Evaluating Against the Test Set. Upon applying the revision, several actions are triggered automatically. First, the Prompt Instruction panel is updated with the new version. Second, the original failed example and the additional examples from the neighborhood probe are added to the Test Set panel. Third, the system runs the new prompt instruction against all test cases.
@@ 4.4. Implementation Details
Our system is implemented as a web-based application that runs directly in the browser. The user interface is built with React, while components that require LLM calls are handled by a Python server developed with Flask. The server communicates with external LLM APIs to support various stages of the workflow.
@@
We primarily used Gemini 2.5 Flash (Comanici et al., 2025) from Google Gemini API [^2] https://ai.google.dev/gemini-api/docs/models to power the workflow steps, such as generating synthetic examples, retrieving similar examples, and suggesting prompt revisions. To ensure fast interactions, reasoning-heavy “thinking mode” was avoided in most cases. For model responses to end users, where prompt instructions are simply combined with user inputs, we employed lighter models such as Gemini 2.5 Flash Lite.
Because the LLM-as-judge module runs at every iteration, its runtime must remain low as the test set grows. To ensure responsiveness, we parallelize these LLM calls. A small scalability experiment shows a near-linear speedup. Evaluating 50 examples decreased elapsed time from 52.8 seconds to 5.6 seconds with 10 workers (9.4x speedup) and to 3.4 seconds with 20 workers (15.6x speedup). Although latency is bounded by API response time, parallel processing effectively mitigates this bottleneck and keeps interaction real-time even for larger test sets.
5. Technical Experiments
We conducted simulation-based experiments to evaluate the workflow’s core components: failure discovery and instruction refinement. The simulation serves two purposes: testing how well the workflow adapts to diverse, application-specific policy definitions, and validating components that rely on probabilistic LLM outputs in a controlled setting. To do this, we created deterministic “personas” that act as ground-truth oracles.
5.1. Setup
We utilized two content moderation domains from our user study: an online student forum and restaurant customer reviews (see Section 6.2 for details). To enable deterministic evaluation, we developed a structured taxonomy and a set of personas with fixed policy preferences.
Categorization. We defined a taxonomy of 24 categorical values spanning four dimensions, designed to capture the factors that shape moderation decisions:
  • Severity: emotional intensity of negativity (e.g., neutral, mild, moderate, strong).
  • Target: who or what is being criticized (e.g., individual, group, general public, service).
  • Content category: the primary issue raised (e.g., facilities, food quality, staff behavior).
  • Problematic style: stylistic markers that escalate perceived harm (e.g., profanity, harsh tone, sarcasm, slang).
To map user inputs to these categories, we used a classifier LLM (Gemini 2.5 Flash), which achieved 89% accuracy against human-labeled ground truth. Note that this is the only LLM involved in the experiment pipeline; all moderation decisions made by the personas are fully deterministic, allowing us to evaluate without relying on LLM-based judgments.
Simulated Personas (Oracles). For each domain, we defined four distinct personas governed by deterministic logic (if-then rules) based on the taxonomy values. These rules produce both a binary label (i.e., whether user inputs are problematic) and a corresponding rationale. To ensure consistency, each persona uses a predefined rationale template that converts the triggered rule into a short, human-readable explanation. This simulates the “human rationale” phase of our workflow and provides structured input to the instruction-refinement module. While prior research has used LLMs to simulate user judgment, we intentionally avoided LLM-based user simulators to prevent the risk of circular validation, given that our workflow already relies on LLMs in other components.
  • Persona A (Civility): prioritizes polite wording; rejects any input containing profanity or insults.
  • Persona B (Expressiveness): prioritizes free speech; tolerates harsh tone and slang while rejecting extreme hate speech or threats.
  • Persona C (Fairness): prioritizes fairness to individuals and groups; tolerates strong criticism of situations but flags context-free slurs, sweeping denigration, or calls for harm.
  • Persona D (Constructiveness): emphasizes informational content; accepts strongly worded complaints but flags vague or purely venting.
5.2. Experiment 1: Validity and Coverage of Failure Discovery
The workflow begins by synthesizing “hard” cases. If these cases are trivial or irrelevant, the co-evolution loop cannot meaningfully proceed. We therefore evaluated whether the failure discovery module (Step 1 of the workflow) generates cases that are both valid failures and semantically diverse.
Protocol. For each persona, we generated five failure candidates using our workflow and compared them against two baselines: (1) simple generation, which prompts the model for typical user inputs; and (2) adversarial generation, which prompts the model to generate inputs that cause model failure (similar to red-teaming). We measured the following two metrics, each averaged over 10 runs:
  • Validity: the proportion of generated examples that were judged incorrect according to the persona’s pre-defined determination rules.
  • Coverage: the proportion of generated examples that triggered failures in at least one of the four personas. This indicates that the example lies in a policy grey area where reasonable policies can disagree, rather than representing a trivial violation that every persona would reject in the same way.
Results. Our system outperformed both baselines. For validity, our system achieved a failure discovery rate of $0.635$, performing comparably to the adversarial baseline ($0.497$) and far exceeding simple generation ($0.031$). For coverage, while the adversarial baseline achieved a score of $0.677$, our system achieved $0.917$. These results suggest that the workflow produces nuanced, realistic edge cases that span multiple interpretations of policy, rather than just exploiting model weaknesses.
5.3. Experiment 2: Effectiveness of Instruction Refinement
We next evaluated the instruction refinement module through two questions: (1) Do the workflow components contribute to instruction improvement? and (2) Does the workflow support cumulative improvement across multiple refinement cycles?
5.3.1. Ablation Study
Protocol. We isolated the impact of the workflow’s unique components by comparing our full workflow against a simplified baseline that skips the rationale-generation and neighborhood-verification steps.
Results. The results demonstrate the value of these components. Starting from an initial accuracy of 0.403 on the holdout set (20 examples written by real users collected from the web; details in Section 6.6), the ablated baseline improved performance marginally to 0.428. In contrast, our full workflow reached 0.502 within a single iteration. This suggests that rationale-driven updates generalize more effectively than narrow, case-specific fixes.
5.3.2. Sequential Iterations
Protocol. We examined whether the workflow fosters continuous improvement by simulating a session with eight consecutive cycles of discovery and refinement. Each configuration (two domains x four personas) was run ten times.
Results. Performance improved steadily across iterations. Accuracy increased sharply after the first iteration (from 0.408 to 0.517), followed by consistent incremental gains across subsequent steps: 0.528, 0.554, 0.570, 0.576, and 0.583 (steps 2–6). Performance peaked at 0.591 at step 7 and stabilized at 0.588 by step 8. While performance improves steadily, it begins to plateau in later stages. This plateau likely reflects the fact that our simulation relies on personas that encode deterministic and well-scoped policies, allowing many straightforward issues to be addressed early. The remaining errors tend to lie near subtle policy boundaries, leaving less room for large accuracy gains. Meanwhile, real users may iteratively develop and adjust their policies over time, a dynamic that is not fully captured by our fixed personas.
These experiments demonstrate that the system’s two core mechanisms—failure discovery and instruction refinement—operate as intended in a controlled setting. At the same time, the observed plateau highlights a broader limitation of simulation-based evaluation: real-world policy design is rarely fixed or deterministic, and instead involves iterative sensemaking as human intent evolves. Thus, it is necessary to examine whether these capabilities translate into practical utility for actual human users. We therefore turn to the next section to present a user study comparing our workflow against a human baseline to assess whether it more effectively supports human creativity and iteration.